Find the sum of the arithmetic series using a formula. 2 + 4 + 6 + ... + 374
1
Expert's answer
2012-12-13T08:45:08-0500
Sn=(a1+an)*n/2
in our case a1=2, an=374, n=187
Sn=(2+374)*187/2=(1+187)*187=188*187=35156